A reactant decomposes with a half-life of 157 s when its initial concentration is 0.240 M. When the initial concentration is 0.782 M, this same reactant decomposes with the same half-life of 157 s.If the reaction order is 1 what is the k value ?
Answer #1

(t1/2)1   = 157s             a1    = 0.240M

(t1/2)2   = 157s              a2   = 0.782M

(t1/2)/(t1/2)2   = (a2/a1)^n-1

157/157           = (0.782/0.24)^n-1

1                    = (3.2584)^n-1

(3.2584)^0    = (3.2584)^n-1

n -1   = 0

n = 1

first order reaction

K   = 0.693/t1/2

         = 0.693/157   = 0.004414s-1 >>>>answer

> Where'd you get the 0.693? You did a little tangent with an n=1, i dont understand how that effected leading to your "0.693" out of NOWHERE

Loki Okami Sun, Nov 7, 2021 4:34 AM